OpenStudy (anonymous):

elimination method 7r-6s=36 6r+7s=43 ordered pair solution (-13,6) is this right??

hero (hero):

7r-6s=36 6r+7s=43 r = (36 + 6s)/7 r = (43-7s)/6 r = r 6(36 + 6s) = 7(43 - 7s) 216 + 36s = 301 - 49s 301 - 216 = 85s 85 = 85s 85/85 = s 1 = s

hero (hero):

r = 6

OpenStudy (anonymous):

(7r-6s=36)-6 (6r-7s=43)7 -42r+36s=-216 42r-49s=301 13s=85

OpenStudy (anonymous):

s=85/13

hero (hero):

Chris, (r,s) = (6,1) is the solution

OpenStudy (anonymous):

-85/13 oops

OpenStudy (anonymous):

ok how is what you did elimination method?

hero (hero):

If you want me to prove it to you, I will

hero (hero):

Because I eliminated r variable, then solved for s

hero (hero):

It's just not elimination method you're used to. Besides, my methods get the correct answer. I can't really speak for the elimination methods you're used to.

OpenStudy (anonymous):

please do i have never seen your way

hero (hero):

I invented my way, that's why. If you use x and y as variables you'll see why.

OpenStudy (anonymous):

always inte

hero (hero):

I invented a plethora of approaches to math that make much more sense than methods taught in school. Some of which are pretty silly and potentially confusing.

hero (hero):

You can prove it to yourself by plugging in r = 6 and s = 1 back into the original equations

OpenStudy (anonymous):

Chris Chris Chris..

OpenStudy (anonymous):

lol i asked him to teach me? what?

OpenStudy (anonymous):

i am sure i made a mistake somewhere but i was interested in his method

hero (hero):

The method I use, you've been taught it already.

hero (hero):

y = mx+b If you have an equation with x and y on the same side, and you want to isolate y, say, for graphing reasons, you can do that. In this case, I isolated r, which would be y. I do this so that it's easier to solve systems of equations, also I can graph the equations I isolated on a graph to check them. So I'm doing two thing by using my method. I'm providing a way to check my answer graphically, and I'm finding an easier, simplier way to solve the more difficult systems of equation problems.

OpenStudy (anonymous):

oh i already looked at what you did but to me its a lot like the susbstitution methos instead of the elimination method asked in the question

hero (hero):

7y-6x=36 6y+7x=43 y = (36 + 6x)/7 <---I graph this to check y = (43-7x)/6 <---and this too y = y 6(36 + 6x) = 7(43 - 7x) 216 + 36x = 301 - 49x 301 - 216 = 85x 85 = 85x 85/85 = x 1 = x 6 = y

OpenStudy (anonymous):

you must realize that some instructors would want to see that problem done by instruction instead of substituting for r

OpenStudy (anonymous):

but im not arguing, just making an observation

hero (hero):

I would argue that I did eliminate one variable while solving.

OpenStudy (anonymous):

what you did was solve each equation for r then set them equal

hero (hero):

Either way, no one has yet to post the solution you're looking for.

hero (hero):

I figure that maybe you would but instead you want to argue with me.

OpenStudy (anonymous):

well im justsaying what you did is almost the exact definition of the substitution method

hero (hero):

Chris, everyone is waiting for you to post the "correct" solution.

OpenStudy (anonymous):

well ok ill show you

hero (hero):

I posted my solution because I don't care much for the elimination method as you know it.

OpenStudy (anonymous):

but im not arguing

hero (hero):

Whatever you say man :P

OpenStudy (anonymous):

i unsderstand that, and if the question didnt ask for a specific method i would agree with you

OpenStudy (anonymous):

im not trying to argue with you at all

OpenStudy (anonymous):

(7r-6s=36)-6 (6r-7s=43)7 -42r+36s=-216 42r+49s=301 85s=85 s=1

hero (hero):

I posted a solution since no one had posted one yet. And by technicality, I did indeed eliminate r variable by substitution.

OpenStudy (anonymous):

that would be elimination method

OpenStudy (anonymous):

same answer just different method

hero (hero):

Well, it's about time...only took about a whole hour for someone to do it.

OpenStudy (anonymous):

if they had asked them to do it susbstitution method, you answer is absolutely correct

OpenStudy (anonymous):

there is just two different methods

hero (hero):

If it has to take an hour to post the solution, I'd rather stick with my methods...

OpenStudy (anonymous):

well just realize there are two methods and when it asks you do do it that way for college per say, then the other method would be wrong

OpenStudy (anonymous):

even though the answer is the same

hero (hero):

I'm just happy the correct solution is finally posted. I just posted my solution for fun.

OpenStudy (anonymous):

heh no problem man, just do not want you to think I am arguing

hero (hero):

I know, you're just trying to get your point across while I'm trying to get mine across. No one in their right minds would consider that arguing :P

OpenStudy (anonymous):

lol i call it debating

OpenStudy (anonymous):

one of us will come up with the right answer eventually

hero (hero):

debating = formal argument

OpenStudy (anonymous):

it sounds nicer though

hero (hero):

This must have been how it was like when Leibnitz and Newton was arguing over who invented Calculus

OpenStudy (anonymous):

lol

hero (hero):

Good night Chris...

OpenStudy (anonymous):

night